OpenStudy (anonymous):

Using the distance formula to find the distance between these point (-5,7) And (2,3) ,, round to the nearest tenth?

OpenStudy (anonymous):

Use \(\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}.\)

OpenStudy (anonymous):

I left my calculator & i cant do it with out -_-

OpenStudy (anonymous):

It's easy to get the expression \((x_2-x_1)^2+(y_2-y_1)^2\) simplified. Once you do that, just google \(\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\). (You have to type sqrt({put the number here}))

OpenStudy (anonymous):

(2-5)^2 + (3+7)^2

OpenStudy (anonymous):

i put this in google?

OpenStudy (anonymous):

@Limitless

OpenStudy (anonymous):

(2-(-5))^2=(2+5)^2 (3-(7))^2=(3-7)^2 Type sqrt((2+5)^2+(3-7)^2) into google.

OpenStudy (anonymous):

8.06225775 is what they come up to . one of my answer choices is 8.7 . would tat be correct?

OpenStudy (anonymous):

or 8.1 @Limitless

OpenStudy (anonymous):

\(8.1\). You round to the nearest tenth. Since the hundredth's place is \(6\), you round up and get \(8.1.\)

OpenStudy (anonymous):

thanks ! ( ;